# Yongtao Li, Lihua Feng, Yuejian Peng

### v1

@@ 3 Preliminaries
The inequality can be written as the following versions:
@@
This inequality was initially obtained by Bollobás and Nikiforov [7, Theorem 1], and it was independently discovered by Cioabă, Feng, Tait and Zhang [9]. The case of equality was recently characterized by Ning and Zhai [45].
@@
We refer to [4, p. 297] and [33, p. 443] for detailed proofs. This result gives the supersaturation on triangles for graphs with more than $n^{2}/4$ edges. Next, we present a generalization for a graph with less than $n^{2}/4$ edges whenever this graph is far from being bipartite. More precisely, for $\varepsilon>0$, we say that a graph $G$ is $\varepsilon$-far from being bipartite if $G^{\prime}$ is not bipartite for every subgraph $G^{\prime}\subseteq G$ with $e(G^{\prime})>e(G)-\varepsilon$. In other words, if $G$ is $\varepsilon$-far from being bipartite, then we need to delete at least $\varepsilon$ edges from $G$ to make it bipartite.
@@
The supersaturation on triangles is usually deduced from Moon–Moser’s inequality. In fact, Theorem 3.3 can also imply the supersaturation on triangles. Note that every bipartite graph of order $n$ has at most $n^{2}/4$ edges. Thus, if $G$ has at least ${n^{2}}/{4}+q$ edges, then it is $q$-far from being bipartite. Theorem 3.3 implies that $G$ contains at least ${qn}/{3}$ triangles. In particular, setting $q=\varepsilon n^{2}$, we obtain that $G$ has at least ${\varepsilon n^{3}}/{3}$ triangles.
For dealing with some extremal graph problems, it is a great advantage that applying Theorem 3.3 can help us to get rid of the use of triangle removal lemma so that we do not require the order of a graph to be sufficiently large. This approach could be considered to be a key innovation in our paper, and it was proven to be effective and feasible in many extremal graph problems. We refer to [26, 27] for more related applications. In what follows, we provide a warm-up practice that has its independent interest. We concentrate on explaining the main ideas which we believe have wide applicability.
@@
First of all, if $G$ is a complete bipartite graph, then $e(G)\leq\lfloor n^{2}/4\rfloor$. Moreover, we have $e(G)=\lambda^{2}\geq n^{2}/4$. It follows that $n$ is even and $G=K_{\frac{n}{2},\frac{n}{2}}$. In this case, the desired result holds immediately. Now, we consider the case that $G$ is not a complete bipartite graph, then the inequality in Lemma 3.2 holds strictly. So we get $e(G)>\lambda^{2}-(3t)/\lambda\geq\lambda^{2}-(6t)/n\geq n^{2}/4-3k$. Theorem 3.3 implies that $G$ is not $6k$-far from being bipartite. Otherwise, if $G$ is $6k$-far from being bipartite, then $t(G)\geq{n}/{6}\cdot(e(G)+6k-n^{2}/4)>kn/2$, which contradicts with the assumption. Hence $G$ is not $6k$-far from being bipartite. Then there exists a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6k$. Moreover, we can get $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4-9k$. Consequently, we obtain that $n/2-3\sqrt{k}<|S|,|T|<n/2+3\sqrt{k}$. Note that $e(K_{|S|,|T|})-e(S,T)<|S||T|-(n^{2}/4-9k)\leq 9k$. In other words, there are less than $9k$ missing edges between $S$ and $T$. Consequently, we get $\delta(G)>(n/2-3\sqrt{k})-9k\geq n/2-12k$. Note that $e(S)+e(T)<6k$. Then it follows that $\Delta(G)<(n/2+3\sqrt{k})+6k\leq n/2+9k$, as desired. ∎
@@
Note that the case $q=1$ in Problem 2.1 was proved by Li, Lu and Peng [25]. Next, we shall consider the case $q\geq 2$ and $n\geq 300q^{2}$ although our proof works for $q=1$. To avoid unnecessary computations, we did not make much efforts to optimize the bound on the order of a graph. Throughout this section, let $G$ be a graph on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$. Suppose on the contrary that $G$ contains less than $q\lfloor{n}/{2}\rfloor$ triangles. In the sequel, we will deduce a contradiction. The essential observation in our proof attributes to the supersaturation-stability as described in Theorem 3.3.
To start the proof of Theorem 2.2, we outline the main steps as follows.
  • Firstly, we establish a bipartition of vertices of $G$. Using Lemma 3.2, we get that $G$ has more than $n^{2}/4-3q$ edges. The supersaturation-stability in Theorem 3.3 implies that one can delete $6q$ edges from $G$ to make it bipartite. So we get a partition $V(G)=S\cup T$ with $e(S)+e(T)<6q$ and $|S|,|T|\approx n/2\pm 3\sqrt{q}$.
  • Secondly, we shall show further that $e(S)+e(T)\leq q$. Moreover, the minimum degree of $G$ is at least $n/2-4q$; see Lemmas 4.2 and 4.3. Let $\mathbf{x}\in\mathbb{R}^{n}$ be the Perron vector of $G$ satisfying $\max\{x_{v}:v\in V(G)\}=1$. Then we will prove in Lemma 4.4 that ${x}_{u}\geq 1-{30q}/{n}$ for every $u\in V(G)$.
  • Finally, we prove in Lemma 4.6 that the bipartition of $G$ is balanced, i.e., $\bigl{|}|S|-|T|\bigr{|}\leq 1$. Using the above lemmas, we shall prove that $G$ has at least $q\lfloor n/2\rfloor$ triangles, or $G$ satisfies $\lambda(G)<\lambda(Y_{n,2,q})$. This leads to a contridiction.
@@
$\frac{n}{2}-3\sqrt{q}<|S|,|T|<\frac{n}{2}+3\sqrt{q}.$
@@
First of all, we know from Lemma 3.1 that $\lambda^{2}(Y_{n,2,1})>\lfloor n^{2}/4\rfloor+2$. Then $\lambda^{2}(G)\geq\lambda^{2}(Y_{n,2,q})>\lfloor n^{2}/4\rfloor+2$. Since $t(G)<q\lfloor{n}/{2}\rfloor$, by Lemma 3.2, we have
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n}>\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q.$ (1)
We claim that $G$ is not $6q$-far from being bipartite. Otherwise, if $G$ is $6q$-far from being bipartite, then using Theorem 3.3, we get
$t(G)\geq\frac{n}{6}\left(e(G)+6q-\frac{n^{2}}{4}\right)>q\left\lfloor\frac{n}{2}\right\rfloor,$
which contradicts with the assumption. So $G$ is not $6q$-far from being bipartite. That is to say, we can delete less than $6q$ edges from $G$ to make it bipartite, i.e., there exists a vertex partition $V(G)=S\cup T$ such that
$e(S)+e(T)<6q.$
@@
$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-9q.$
Without loss of generality, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq{n}/{2}-3\sqrt{q}$. Then $|T|=n-|S|\geq{n}/{2}+3\sqrt{q}$. It yields that
$e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-3\sqrt{q}\right)\left(\frac{n}{2}+3\sqrt{q}\right)=\frac{n^{2}}{4}-9q,$
a contradiction. Thus, we must have ${n}/{2}-3\sqrt{q}<|S|,|T|<{n}/{2}+3\sqrt{q}$. ∎
@@
Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$. Suppose on the contrary that $e(S)+e(T)\geq q+1$. We write $K_{|S|,|T|}$ for the complete bipartite graph on the vertex parts $S$ and $T$. Note that each edge of $G[S]$ or $G[T]$ can yield at least $\min\{|S|,|T|\}$ triangles in $K_{|S|,|T|}$. Recall in Lemma 4.1 that $e(S)+e(T)<6q$. So each missing edge between $S$ and $T$ destroys at most $6q$ triangles. By Lemma 4.1 again, we have $\min\{|S|,|T|\}>{n}/{2}-3\sqrt{q}$, and $G[S,T]$ misses at most $9q$ edges from the graph $K_{|S|,|T|}$. Hence, we conclude
$t(G)\geq(q+1)\left(\frac{n}{2}-3\sqrt{q}\right)-9q\cdot 6q>(q+1)\frac{n}{2}-58q^{2}\geq q\left\lfloor\frac{n}{2}\right\rfloor,$
@@
We have
$e(S,T)>\frac{n^{2}}{4}+2-4q$
@@
$\frac{n}{2}-2\sqrt{q}<|S|,|T|<\frac{n}{2}+2\sqrt{q}.$
@@
$\frac{n}{2}-4q\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+q+2\sqrt{q}.$
@@
By Lemma 4.2, we obtain $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4+2-4q$. By the AM-GM inequality, we can get ${n}/{2}-2\sqrt{q}\leq|S|,|T|\leq{n}/{2}+2\sqrt{q}$.
By Rayleigh’s formula, it is well-known that $\delta(G)\leq\lambda(G)\leq\Delta(G)$. Since $e(S)+e(T)\leq q$, each vertex of $S$ has degree at most $q$ in $G[S]$, and this also holds for the vertices of $T$. Then it follows that $\Delta(G)\leq\max\{|S|,|T|\}+q\leq{n}/{2}+2\sqrt{q}+q$. Next, we show that $\delta(G)\geq{n}/{2}-4q$. Otherwise, if there exists a vertex $v\in V(G)$ with the minimum degree and $d(v)\leq{n}/{2}-4q-{1}/{2}$, then
$e(G\setminus\{v\})=e(G)-d(v)$$\geq$ $\left\lfloor\frac{n^{2}}{4}\right\rfloor+2-3q-\left(\frac{n}{2}-4q-\frac{1}{2}\right)$
$\geq$ $\frac{(n-1)^{2}}{4}+q+2.$
By Theorem 1.2, we can find $(q+2)\lfloor\frac{n-1}{2}\rfloor$ triangles in $G\setminus\{v\}$. This is a contradiction with the assumption $t(G)<q\lfloor{n}/{2}\rfloor$. Thus, we must have $\delta(G)\geq{n}/{2}-4q$. ∎
Recall that $\mathbf{x}\in\mathbb{R}^{n}$ is the Perron vector of $G$ with $\max\{x_{v}:v\in V(G)\}=1$.
@@
Let $z\in V(G)$ be a vertex such that ${x}_{z}$ attains the maximum eigenvector entry of $\mathbf{x}$ and ${x}_{z}=1$. Without loss of generality, we may assume that $z\in S$. Our proof of this lemma falls naturally into two steps.
Step 1. For every $u\in S$, we have ${x}_{u}>1-\frac{15q}{n}$.
Since $e(S)\leq q$ by Lemma 4.2, we get $d_{S}(u)\leq q$. By Lemma 4.3, we have $d_{T}(u)=d(u)-d_{S}(u)\geq({n}/{2}-4q)-q={n}/{2}-5q$. By Lemma 4.3 again, we get $|T|\leq{n}/{2}+2\sqrt{q}$. Then
@@
$\leq$ $-\left(\frac{n}{2}-5q\right)+\left(\frac{n}{2}+2\sqrt{q}\right)$
$=$ $5q+2\sqrt{q}.$
@@
$\lambda{x}_{u}-\lambda{x}_{z}\geq-\Bigl{(}|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|\Bigr{)}-|N_{S}(z)|\geq-6q-2\sqrt{q}.$
@@
${x}_{u}\geq 1-\frac{6q+2\sqrt{q}}{\lambda}>1-\frac{12q+4\sqrt{q}}{n}>1-\frac{15q}{n}.$
Step 2. For every $u\in T$, we have ${x}_{u}>1-\frac{30q}{n}$.
@@
$\lambda{x}_{u}=\sum_{v\in N(u)}{x}_{v}\geq\sum_{v\in N_{S}(u)}{x}_{v}\geq\left(1-\frac{12q+4\sqrt{q}}{n}\right)d_{S}(u).$
By Lemma 4.3, we can see that $d(u)\geq\delta(G)\geq{n}/{2}-4q$, which together with $d_{T}(u)\leq q$ gives $d_{S}(u)=d(u)-d_{T}(u)\geq{n}/{2}-5q$. Hence, for $q\geq 2$, we have
${x}_{u}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})d_{S}(u)}{\lambda}\geq\frac{(1-\frac{12q+4\sqrt{q}}{n})(\frac{n}{2}-5q)}{\frac{n}{2}+q+2\sqrt{q}}>1-\frac{30q}{n}.$
@@
Using Lemma 4.4, namely, ${x}_{u}=1-O({1}/{n})$, we will show that the bipartition $V(G)=S\cup T$ is balanced (see the forthcoming Lemma 4.6). To start with, we fix some notation for convenience. Let $K_{|S|,|T|}$ be the complete bipartite graph with partite sets $S$ and $T$. Let $G_{1}=G[S]\cup G[T]$ be the graph whose edges consist of the class-edges, and let $G_{2}$ be the graph on $V(G)$ with the missing edges between $S$ and $T$, that is, $E(G_{2})=E(K_{|S|,|T|})\setminus E(G)$. Note that $e(G)=e(G_{1})+e(K_{|S|,|T|})-e(G_{2})$.
@@
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}\leq\frac{120q^{2}}{n(n-60q)}.$
@@
$\mathbf{x}^{\top}\mathbf{x}>n\left(1-\frac{30q}{n}\right)^{2}>n\left(1-\frac{60q}{n}\right)=n-60q.$
@@
$\lambda(G)$ $=$ $\frac{\mathbf{x}^{\top}\bigl{(}A(K_{|S|,|T|})+A(G_{1})-A(G_{2})\bigr{)}\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
$=$ $\frac{\mathbf{x}^{\top}A(K_{|S|,|T|})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}+\frac{\mathbf{x}^{\top}A(G_{1})\mathbf{x}}{\mathbf{x}^{\top}\mathbf{x}}$
$<$ $\sqrt{|S||T|}+\frac{2q}{n-60q}.$
@@
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<2q\left(\frac{1}{n-60q}-\frac{1}{n}\right)=\frac{120q^{2}}{n(n-60q)}.$
@@
$\frac{1}{n}<\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{120q^{2}}{n(n-60q)}.$
@@
By Lemma 4.6, we know that $\bigl{|}|S|-|T|\bigr{|}\leq 1$. Recall that $Y_{n,2,q}$ is the graph on the vertex sets $S$ and $T$ with all edges between $S$ and $T$, together with $q$ pairwise disjoint edges embedded into the vertex part of size $\lceil{n}/{2}\rceil$. Clearly, we have $e(Y_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ edges. We denote

### v2

@@ 3.2 Counting the number of triangles
As a consequence, Lemma 3.4 implies an interesting result of Nosal [50], which states that every $m$-edge graph $G$ with $\lambda(G)>\sqrt{m}$ contains a triangle; see [41, 47] for related results. The inequality in Lemma 3.4 can be rephrased as the following versions:
@@
This inequality was initially obtained by Bollobás and Nikiforov [7, Theorem 1], and it was independently discovered by Cioabă, Feng, Tait and Zhang [9], as well as Ning and Zhai [48]. The case of equality was characterized by Ning and Zhai [48].
@@
We refer to [4, p. 297] and [36, p. 443] for detailed proofs. This result gives the supersaturation on triangles for graphs with more than $n^{2}/4$ edges. Next, we present a generalization for a graph with less than $n^{2}/4$ edges whenever this graph is far from being bipartite. More precisely, for $\varepsilon>0$, we say that a graph $G$ is $\varepsilon$-far from being bipartite if any subgraph $G^{\prime}$ of $G$ with more than $e(G)-\varepsilon$ edges is not bipartite. In other words, if $G$ is $\varepsilon$-far from being bipartite, then we must delete at least $\varepsilon$ edges from $G$ to make it bipartite.
@@
For the convenience of the reader, we include a proof here.
Proof.
For each vertex $v\in V(G)$, let $N_{v}=N(v)$ and $N_{v}^{c}=V(G)\setminus N(v)$. Since $G$ is $t$-far from being bipartite, we must have $e(N_{v})+e(N_{v}^{c})\geq\varepsilon$ for every $v\in V(G)$. Indeed, if $e(N_{v})+e(N_{v}^{c})<\varepsilon$ for some $v$, then deleting all edges inside $N_{v}$ and $N_{v}^{c}$ yields a bipartite subgraph with more than $e(G)-\varepsilon$ edges, a contradiction. For each $v\in V(G)$, we have
$\sum_{w\in N_{v}^{c}}d(w)=2e(N_{v}^{c})+e(N_{v}^{c},N_{v})=e(N_{v}^{c})+m-e(N_{v})\geq m+\varepsilon-2e(N_{v}).$
Summing this inequality over all vertices $v\in V(G)$ gives
$\sum_{v\in V}\sum_{w\in N_{v}^{c}}d(w)\geq mn+\varepsilon n-2\sum_{v\in V}e(N_{v})=mn+\varepsilon n-6t(G),$
where we used the identity $3t(G)=\sum_{v\in V}e(N_{v})$. On the other hand, we get
$\sum_{v\in V}\sum_{w\in N_{v}^{c}}d(w)=\sum_{v\in V}\left(2m-\sum_{w\in N_{v}}d(w)\right)=2mn-\sum_{w\in V}d^{2}(w).$
Combining the two inequalities, we obtain
$6t(G)\geq\varepsilon n-mn+\sum_{w\in V}d^{2}(w)\geq\varepsilon n-mn+\frac{4m^{2}}{n},$
where the last inequality follows by Cauchy–Schwarz’s inequality. Note that ${4m^{2}}/{n}\geq 2mn-{n^{3}}/{4}$. Substituting this into the previous inequality gives the desired bound. ∎
The supersaturation on triangles is usually deduced from Moon–Moser’s inequality. In fact, Theorem 3.5 can also imply the supersaturation on triangles. Note that every bipartite graph of order $n$ has at most $n^{2}/4$ edges. Thus, if $G$ has at least ${n^{2}}/{4}+q$ edges, then it is $q$-far from being bipartite. Theorem 3.5 implies that $G$ contains at least qn/3{qn}/{3} triangles. In particular, setting $q=\varepsilon n^{2}$, we obtain that $G$ has at least ${\varepsilon n^{3}}/{3}$ triangles.
For dealing with some extremal graph problems, it is a great advantage that applying Theorem 3.5 can help us to get rid of the use of triangle removal lemma so that we do not require the order of a graph to be sufficiently large. This approach could be considered to be a key ingredient in our paper, and it was proven to be effective and feasible in many extremal graph problems. For other related applications, we refer the reader to [28, 29].
3.3 A spectral stability for graphs with few triangles
In what follows, we provide a warm-up practice that has its independent interest. We concentrate on explaining the main ideas which we believe have wide applicability.
@@ 3.3 A spectral stability for graphs with few triangles
First, if $G$ is a complete bipartite graph, then $e(G)\leq\lfloor n^{2}/4\rfloor$. Since $e(G)=\lambda^{2}(G)\geq n^{2}/4$, it follows that $n$ is even and $G=K_{\frac{n}{2},\frac{n}{2}}$. In this case, the desired result holds immediately. Now suppose that $G$ is not a complete bipartite graph. Then the inequality in Lemma 3.4 is strict. So we get $e(G)>\lambda^{2}-(3t)/\lambda\geq\lambda^{2}-(6t)/n\geq n^{2}/4-3k$. Theorem 3.5 implies that $G$ is not $6k$-far from being bipartite. Otherwise, if $G$ is $6k$-far from being bipartite, then $t(G)\geq{n}/{6}\cdot(e(G)+6k-n^{2}/4)>kn/2$, which contradicts the assumption. Hence $G$ is not $6k$-far from being bipartite. Consequently, there exists a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6k$. Moreover, we get $e(S,T)=e(G)-e(S)-e(T)>n^{2}/4-9k$. It follows that $n/2-3\sqrt{k}<|S|,|T|<n/2+3\sqrt{k}$. Note that $e(K_{|S|,|T|})-e(S,T)<|S||T|-(n^{2}/4-9k)\leq 9k$. In other words, there are fewer than $9k$ missing edges between $S$ and $T$. Consequently, we get $\delta(G)>(n/2-3\sqrt{k})-9k\geq n/2-12k$. Finally, since $e(S)+e(T)<6k$, we have $\Delta(G)<(n/2+3\sqrt{k})+6k\leq n/2+9k$, as desired. ∎
@@
The case $q=1$ was proved by Li, Lu and Peng [27]. Next, we consider the case $2\leq q\leq\frac{1}{11}\sqrt{n}$ although our proof works for $q=1$. Throughout this section, let $G$ be a graph on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$. Suppose on the contrary that $t(G)<q\lfloor{n}/{2}\rfloor$. In the sequel, we will deduce a contradiction. The key insight of our proof lies in the application of the supersaturation-stability (Theorem 3.5). We outline the main steps as follows.
  • Using Lemma 3.4, we get that $G$ contains more than $\lfloor n^{2}/4\rfloor-q$ edges. Theorem 3.5 implies that we can delete less than $4q$ edges from $G$ to make it bipartite. So we get a bipartition $V(G)=S\cup T$ with $e(S)+e(T)<4q$ and $|S|,|T|\approx\frac{n}{2}\pm\sqrt{5q}$.
  • Secondly, we show that $e(S)+e(T)\leq q$ and $\delta(G)\geq\frac{n}{2}-2q$; see Lemmas 4.2 and 4.3. Let $\bm{x}\in\mathbb{R}^{n}$ be the Perron vector of $G$ satisfying $\max\{x_{v}:v\in V(G)\}=1$. Then we will prove in Lemma 4.4 that ${x}_{u}\geq 1-{20q}/{n}$ for every $u\in V(G)$.
  • We prove in Lemma 4.6 that the bipartition of $G$ is balanced, i.e., $\bigl||S|-|T|\bigr|\leq 1$. Then we shall prove that $t(G)\geq q\lfloor n/2\rfloor$, or $G$ satisfies $\lambda(G)<\lambda(Y_{n,2,q})$, a contridiction.
@@
$\frac{n}{2}-\sqrt{5q}<|S|,|T|<\frac{n}{2}+\sqrt{5q}.$
@@
By Lemma 3.1, we know that $\lambda^{2}(G)\geq\lambda^{2}(Y_{n,2,q})>\lfloor n^{2}/4\rfloor+2q$. Since $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)>\frac{n}{2}$, by Lemma 3.4, we have
$e(G)\geq\lambda^{2}(G)-\frac{3t(G)}{\lambda(G)}>\lambda^{2}(G)-\frac{6t(G)}{n}>\left\lfloor\frac{n^{2}}{4}\right\rfloor-q.$ (1)
We claim that $G$ is not $4q$-far from being bipartite. Otherwise, if $G$ is $4q$-far from being bipartite, then using Theorem 3.5, we get
$t(G)\geq\frac{n}{6}\left(e(G)+4q-\frac{n^{2}}{4}\right)>q\left\lfloor\frac{n}{2}\right\rfloor,$
which contradicts with the assumption. So $G$ is not $4q$-far from being bipartite. That is to say, we can delete less than $4q$ edges from $G$ to make it bipartite, i.e., there exists a vertex partition $V(G)=S\cup T$ such that
$e(S)+e(T)<4q.$
@@
$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-5q.$
Without loss of generality, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq\frac{n}{2}-\sqrt{5q}$. Then $|T|=n-|S|\geq\frac{n}{2}+\sqrt{5q}$. It yields that
$e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-\sqrt{5q}\right)\left(\frac{n}{2}+\sqrt{5q}\right)=\frac{n^{2}}{4}-5q,$
a contradiction. Thus, we must have $\frac{n}{2}-\sqrt{5q}<|S|,|T|<\frac{n}{2}+\sqrt{5q}$. ∎
@@
Suppose on the contrary that $e(S)+e(T)=k$, where $k$ is an integer with $q+1\leq k<4q$ by Lemma 4.1. For simplicity, we write $K_{S,T}$ for the complete bipartite graph on the vertex parts $S$ and $T$. Note that every triangle of $G$ must contain an edge from $G[S]\cup G[T]$, and each edge of $G[S]\cup G[T]$ yields at least $\min\{|S|,|T|\}$ triangles. Each missing edge between $S$ and $T$ destroys at most $k$ triangles. By Lemma 4.1, we have $\min\{|S|,|T|\}>\frac{n}{2}-\sqrt{5q}$, and $G[S,T]$ misses at most $5q$ edges from $K_{S,T}$. We conclude
$t(G)>k\left(\frac{n}{2}-\sqrt{5q}\right)-5q\cdot k\geq(q+1)\left(\frac{n}{2}-\sqrt{5q}-5q\right)\geq q\left\lfloor\frac{n}{2}\right\rfloor,$
@@
With the previous notation, we have
$e(S,T)>\frac{n^{2}}{4}-2q$
@@
$\frac{n}{2}-\sqrt{2q}<|S|,|T|<\frac{n}{2}+\sqrt{2q}.$
@@
$\frac{n}{2}-2q\leq\delta(G)\leq\lambda(G)\leq\Delta(G)\leq\frac{n}{2}+q+\sqrt{2q}.$
@@
Combining (1) with Lemma 4.2, we obtain $e(S,T)=e(G)-e(S)-e(T)>\lfloor n^{2}/4\rfloor-2q$. By the AM-GM inequality, we can get $\frac{n}{2}-\sqrt{2q}\leq|S|,|T|\leq\frac{n}{2}+\sqrt{2q}$.
By the Rayleigh formula, it is well-known that $\delta(G)\leq\lambda(G)\leq\Delta(G)$. Since $e(S)+e(T)\leq q$, each vertex of $S$ has degree at most $q$ in $G[S]$, and this also holds for the vertices of $T$. Then it follows that $\Delta(G)\leq\max\{|S|,|T|\}+q\leq\frac{n}{2}+\sqrt{2q}+q$. Next, we show that $\delta(G)\geq\frac{n}{2}-2q$. Otherwise, if there exists a vertex $v\in V(G)$ with degree and $d(v)\leq\frac{n}{2}-2q-\frac{1}{2}$, then using (1) gives
$e(G\setminus\{v\})=e(G)-d(v)$ $\geq$ $\left(\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-q+1\right)-\left(\frac{n}{2}-2q-\frac{1}{2}\right)$
$\geq$ $\frac{(n-1)^{2}}{4}+q+1.$
By Theorem 1.2, we can find at least $(q+1)\lfloor\frac{n-1}{2}\rfloor$ triangles in $G\setminus\{v\}$. This is a contradiction with the assumption $t(G)<q\lfloor\frac{n}{2}\rfloor$. Thus, we must have $\delta(G)\geq\frac{n}{2}-2q$. ∎
Recall that $\bm{x}\in\mathbb{R}^{n}$ is the Perron vector of $G$ with $\max\{x_{v}:v\in V(G)\}=1$.
@@
Let $z\in V(G)$ be a vertex such that ${x}_{z}$ attains the maximum eigenvector entry of $\bm{x}$ and ${x}_{z}=1$. Without loss of generality, we may assume that $z\in S$. Our proof of this lemma falls naturally into two steps.
Step 1. For every $u\in S$, we have ${x}_{u}>1-\frac{11q}{n}$.
Using Lemma 4.2 gives $e(S)\leq q$ and $d_{S}(u)\leq q$. By Lemma 4.3, we have $d_{T}(u)=d(u)-d_{S}(u)\geq(\frac{n}{2}-2q)-q=\frac{n}{2}-3q$. By Lemma 4.3 again, we get $|T|\leq\frac{n}{2}+\sqrt{2q}$. Then
@@
$\leq$ $-\left(\frac{n}{2}-3q\right)+\left(\frac{n}{2}+\sqrt{2q}\right)$
$=$ $3q+\sqrt{2q}.$
@@
$\lambda{x}_{u}-\lambda{x}_{z}\geq-\Bigl(|N_{T}(z)|-|N_{T}(u)\cap N_{T}(z)|\Bigr)-|N_{S}(z)|\geq-4q-\sqrt{2q}.$
@@
${x}_{u}\geq 1-\frac{4q+\sqrt{2q}}{\lambda}>1-\frac{8q+2\sqrt{2q}}{n}>1-\frac{11q}{n}.$
Step 2. For every $u\in T$, we have ${x}_{u}>1-\frac{20q}{n}$.
@@
$\lambda{x}_{u}=\sum_{v\in N(u)}{x}_{v}\geq\sum_{v\in N_{S}(u)}{x}_{v}>\left(1-\frac{8q+2\sqrt{2q}}{n}\right)d_{S}(u).$
By Lemma 4.3, we can see that $d(u)\geq\delta(G)\geq\frac{n}{2}-2q$, which together with $d_{T}(u)\leq q$ gives $d_{S}(u)=d(u)-d_{T}(u)\geq\frac{n}{2}-3q$. Hence, for $q\geq 2$, we have
${x}_{u}>\frac{(1-\frac{8q+2\sqrt{2q}}{n})d_{S}(u)}{\lambda}\geq\frac{(1-\frac{8q+2\sqrt{2q}}{n})(\frac{n}{2}-3q)}{\frac{n}{2}+q+\sqrt{2q}}\geq 1-\frac{20q}{n}.$
@@
Lemma 4.4 gives ${x}_{u}>1-\frac{20q}{n}$, we will show that the bipartition $V(G)=S\cup T$ is balanced (see Lemma 4.6). To start with, we fix some notation. Let $K_{S,T}$ be the complete bipartite graph with partite sets $S$ and $T$. Let $G_{1}=G[S]\cup G[T]$ be the graph whose edges consist of the class-edges, and let $G_{2}$ be the graph on $V(G)$ with the missing edges between $S$ and $T$, that is, $E(G_{2})=E(K_{S,T})\setminus E(G)$. Note that $e(G)=e(G_{1})+e(K_{S,T})-e(G_{2})$.
@@
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{80q^{2}}{n(n-40q)}.$
@@
$\bm{x}^{\mathrm{T}}\bm{x}>n\left(1-\frac{20q}{n}\right)^{2}>n\left(1-\frac{40q}{n}\right)=n-40q.$
@@
$\lambda(G)$ $=$ $\frac{\bm{x}^{\mathrm{T}}\bigl(A(K_{S,T})+A(G_{1})-A(G_{2})\bigr)\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$
$=$ $\frac{\bm{x}^{\mathrm{T}}A(K_{S,T})\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}+\frac{\bm{x}^{\mathrm{T}}A(G_{1})\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$
$<$ $\sqrt{|S||T|}+\frac{2q}{n-40q}.$
@@
$\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<2q\left(\frac{1}{n-40q}-\frac{1}{n}\right)=\frac{80q^{2}}{n(n-40q)}.$
@@
$\frac{1}{n}<\frac{2}{n}\left\lfloor\frac{n^{2}}{4}\right\rfloor-\sqrt{|S||T|}<\frac{80q^{2}}{n(n-40q)}.$
@@
By Lemma 4.6, we know that $\bigl||S|-|T|\bigr|\leq 1$. Recall that $Y_{n,2,q}$ is the graph on the vertex sets $S$ and $T$ with all edges between $S$ and $T$, together with $q$ pairwise disjoint edges embedded into the larger partite set. Clearly, we have $e(Y_{n,2,q})=\lfloor n^{2}/4\rfloor+q$. We denote